OpenStudy (anonymous):

HELP WILL GIVE MEDALS. Trapezoid JKLM is shown on the coordinate plane below: Trapezoid JKLM on the coordinate plane with ordered pairs at J negative 2, 1, at K 1, 1, at L 3, negative 2, at M negative 4, negative 2. If trapezoid JKLM is translated according to the rule (x, y) → (x + 5, y − 4), what are the coordinates of point L'?

OpenStudy (anonymous):

@imqwerty @CallMeKiki @HELPMEPLZ!! @Buttercup214 @Ghostgate

OpenStudy (anonymous):

Help please! anyone!?

OpenStudy (ghostgate):

Well, if I'm reading this correctly, your first trapezoid has coordinates: J(-2, 1), K(1, 1), L(3, -2), and then M(-4, -2). Now it gives you the transition equation for how the trapezoid would move which is: (x, y) -> (x + 5, y - 4) Since you are looking for L, simply place L's coordinates inside: L = (3 + 5, -2 - 4) - Now solve it. L = (8, -6) - This should be the new set of coordinates for L. I hope this helps in someway! Have a great day! ;) {---Ghostgate---}

imqwerty (imqwerty):

u just add 5 to the x cordinate and subtract 4 from the y cordinate :)

OpenStudy (anonymous):

Oh, okay. this one.. Hexagon DEFGHI is translated on the coordinate plane below to create hexagon D'E'F'G'H'I': Hexagon DEFGHI and Hexagon D prime E prime F prime G prime H prime I prime on the coordinate plane with ordered pairs at D are 3, 5, at E 7, 5, at F 8, 2, at G 7, negative 1, at H 3, negative 1, at I 2, 2; at D prime negative 6, 2, at E prime negative 2, 2, at F prime negative 1, negative 1, at G prime negative 2, negative 4, at H prime negative 6, negative 4, at I prime negative 7, negative 1 Which rule represents the translation of hexagon DEFGHI to hexagon D'E'F'G'H'I'? (x, y)→(x − 9, y − 3) (x, y)→(x − 3, y − 9) (x, y)→(x + 3, y + 3) (x, y)→(x + 9, y + 9) What do i have to do on this? just figure out what to add and subtract to to the same?

OpenStudy (anonymous):

@Ghostgate

imqwerty (imqwerty):

jst consider a few points like i chose D and G :D so D=(3,5) and G=(7,-1) and now we see D' and G' D'=(-6,2) G'=(-2,-4) and now u just have to check the options out jst try out the options for example lets consider the option 1 it says that (x, y)→(x − 9, y − 3) so we get the D' cordinates when we put x and y cordinates of D in (x − 9, y − 3) so we put D=(3,5) in this eq we get- D'=(3-9 , 5-3) D'=(-6,2) :O this one is correct lets check for G' we must get G' cordinates when we put the cordinates of G on (x − 9, y − 3) and we have G=(7,-1) so G'=(7-9, -1-3) G'=(-2,-4) :O this one is also correct and we can see that no other option follows this rule so option 1 is corect

OpenStudy (anonymous):

Thank you guys, I printed this off so i can use it as a reference sheet! You guys helped me alot (:

OpenStudy (ghostgate):

Again, take the various coordinates that each side or letter represents and then add them to the following equation. If it comes out with the prime variables then you have the right equation. DEFGHI has coordinates: D(3, 5), E(7, 5), F(8, 2), G(7, -1), H(3, -1), I(2, 2) D'E'F'G'H'I'(Which are represented as D prime, E prime, F prime, etc.) which has coordinates: D'(-6, 2), E'(-2, 2), F'(-1, -1), G'(-2, -4), H'(-6, -4), I'(-7, -1) Now here are the equations: (x, y)→(x − 9, y − 3) (x, y)→(x − 3, y − 9) (x, y)→(x + 3, y + 3) (x, y)→(x + 9, y + 9) Option 1 is correct due to how each coordinate comes out as the prime coordinate through equation 1. For instance: D(3, 5) -> (3 - 9, 5 - 3) = D' (-6, 2) You can do that for each one and you'll come out with the correct answers. {---Ghostgate---}

OpenStudy (ghostgate):

@lolbug You're welcome! If you need anymore help, just tag me. ;)